OpenStudy (anonymous):

consider this systems of equations 2/3 x +3/5 y =12 5/2 y - 3x=6 the expression that gives the value of x is... the solution for the systems of equations...

OpenStudy (wolf1728):

Do you know how to solve simultaneous equations?

OpenStudy (anonymous):

no

OpenStudy (wolf1728):

A) 2/3 x +3/5 y =12 B) -3x +5/2 y =6 One way is to solve for x in one equation

OpenStudy (anonymous):

I don't get it

OpenStudy (wolf1728):

B) -3x = -5/2 y + 6 We divide both sides by -3 x = 5/6 y -2

OpenStudy (anonymous):

options the expression that gives the value of x is... (5A/3)+(2B/5), (3A/5)-(5B/2), (5A/2)+(3B/5), (5A/3)-(2B/5), (3A/2) + (B/3)

OpenStudy (usukidoll):

mcfizzy STOP! What is that attachment?

OpenStudy (anonymous):

the solution to the systems of equations is (1118/47, 396/47) (33/13, 50/13) (99/13, 150/13) (8,12)

OpenStudy (anonymous):

@UsukiDoll what's wrong?

OpenStudy (usukidoll):

I've seen this poster (mcfizzy) posting an attachment that could be malicious. I've sent a report.

OpenStudy (anonymous):

okay, I probably shouldn't click on the attachment then. I'll send a report too.

OpenStudy (usukidoll):

yeah.. under his profile all of his questions answered has spam with that attachment. everyone please report this.

OpenStudy (anonymous):

I could use help with this problem, I'm still very confused

OpenStudy (wolf1728):

x = 99/13 y = 150/13

OpenStudy (usukidoll):

@wolf1728 is there to help. I have to warn everyone else about this issue as this is serious and against terms and conditions.

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

@wolf1728 I also need the expression that gives the value of x

OpenStudy (wolf1728):

Kaelyn - you do not know how to solve these so why do you need the expression for value of x ?

OpenStudy (anonymous):

because the question asks me to find the expression that finds the value of x, I don't want you to give me the answer just help me find the right one

OpenStudy (wolf1728):

I think you should find someone else to help. I'm quitting for the moment.

OpenStudy (anonymous):

okay thank you for your help

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

May I try?

OpenStudy (usukidoll):

go for it @waterineyes as long as it's not a direct answer

OpenStudy (anonymous):

go ahead @waterineyes

OpenStudy (anonymous):

and @UsukiDoll I agree don't give the answer to me just help me find the answer myself

OpenStudy (anonymous):

Okay, don't let our interaction stop anywhere, just reply if you know or not, just be with me...

OpenStudy (anonymous):

okay

OpenStudy (lynfran):

\[\frac{ 2 }{ 3 }x+\frac{ 3 }{ 4 }y=12\]\[\frac{ 5 }{ 2}y-3x=6\]frm equation 2,\[\frac{ 5 }{ 2 }y=6+3x\]\[y=\frac{ 2(6+3x) }{ 5 }\]now sub y=2(6+3x)/5 in eq 1 \[\frac{ 2 }{ 3 }x+\frac{ 3 }{ 4 }(\frac{ 2(6+3x) }{ 5 })=12\]simplify and solve for x we get \[\frac{ 2 }{ 3 }x+\frac{ 36+18x }{ 20 }=12\]\[\frac{ 40x +108+54x}{ 60 }=12\]\[\frac{ 94x+108 }{ 60 }=12\]multiply both side of the equation we get..\[94x+108=720\]\[94x=720-108\]\[x=612/94\]\[x=6.510638\]...so y =...\[y=\frac{ 2(6+3(6.510638)) }{ 5 }\]\[y=10.21276\]..

OpenStudy (anonymous):

okay understanding so far

OpenStudy (usukidoll):

-_- *throws a water bottle @waterineyes

OpenStudy (anonymous):

Both our equations are in fraction form, we should simplify them to make a good looking equation first... Do you know how to take LCM?

OpenStudy (anonymous):

Hell, my net stopped working.. :(

OpenStudy (anonymous):

Yes I know how to take LCM

OpenStudy (anonymous):

Like in first equation, 15 is the LCM, so you can multiply whole equation by 15...

OpenStudy (anonymous):

@LynFran I am afraid you are right everywhere, but you have taken wrong fraction in very first equation,, 3/5 is good, not 3/4.. if I am not blind then.. :P Your method is good.. :)

OpenStudy (usukidoll):

if we find our y and x, keep it in fraction form. ... decimals are a pain to deal with T_T

OpenStudy (usukidoll):

*throws a canteen at @waterineyes

OpenStudy (anonymous):

Guys lets not throw things...

OpenStudy (lynfran):

o i see im sleepy lol @waterineyes goodnight u guys

OpenStudy (anonymous):

Goodnight @LynFran

OpenStudy (anonymous):

\(10x + 9y = 180\) --- 1 This you got?

OpenStudy (usukidoll):

what is that ?

OpenStudy (usukidoll):

sigh....... the second equation is easier to deal with 2/3 x +3/5 y =12 5/2 y - 3x=6 \[\frac{2}{3}x+\frac{3}{5}y=12\] \[\frac{5}{2}y-3x=6\] so multiply 2 throughout the entire second equation \[\frac{5}{2}y-3x=6\]

OpenStudy (anonymous):

And this way your second equation becomes: \(5y - 6x = 12\) --------- 2

OpenStudy (usukidoll):

uh huh

OpenStudy (anonymous):

@Kaelyn78 are you getting how I formed these two equations?

OpenStudy (anonymous):

yes I understand

OpenStudy (usukidoll):

10x + 9y = 180 5y-6x=12

OpenStudy (anonymous):

Thanks.. Again and again I was going above to see these equations, as you are posting in between so things are going above..

OpenStudy (anonymous):

you can write the second equation, other way round, x first and then y: \(10x + 9y = 180 \\ -6x + 5y = 12\)

OpenStudy (anonymous):

\[[10x + 9y = 180 ] \times 6 = 60x + 45y = 108 \\ [-6x + 5y = 12 ] \times 10 = -60x + 50y = 120\]

OpenStudy (anonymous):

I have multiplied the coefficient of x in second equation with first equation and same for second equation.. are you getting this step?

OpenStudy (anonymous):

yes I understand

OpenStudy (anonymous):

Now you just need to add those equations...

OpenStudy (anonymous):

\[\cancel{60x} + 45y = 1080 \\ \cancel{-60x} + 50y = 120 \\ --------- \\95y = 2200\]

OpenStudy (anonymous):

Sorry, that is 1080 and not 108...

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Getting previous step?

OpenStudy (anonymous):

yes I'm getting it

OpenStudy (anonymous):

Now find y, by dividing by 95 both the sides..

OpenStudy (anonymous):

23?

OpenStudy (anonymous):

Wait, I am just double checking it.. Let me take copy and pen too..

OpenStudy (anonymous):

It is certainly a wrong...

OpenStudy (anonymous):

9 multiplied by 6 is 54 not 45, just keep eyes on what I am doing, I can be wrong anywhere in writing digits. :P

OpenStudy (anonymous):

After Multiplying 6 and 10 the equations you will get are: \(60x + 54y = 1080 \\ -60x + 50y = 120\) Form this you will get: \(104y = 2200\)

OpenStudy (anonymous):

Kill me.....

OpenStudy (anonymous):

1080 + 120 = 1200.. What has happened to me today.. :(

OpenStudy (anonymous):

@waterineyes aw sorry I can't help with that

OpenStudy (anonymous):

\(104y = 1200\) Now just find y by cancelling out the things..

OpenStudy (anonymous):

if it's any consolation I would have gone certifiably insane trying to do this all by myself

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

\[y = \frac{1200}{104}\]

OpenStudy (usukidoll):

gotta reduce that fraction

OpenStudy (anonymous):

Do not use calculator, just cancel it out with 8, you will get in fraction..

OpenStudy (anonymous):

Now, put the fraction in any one of the equation above to find x...

OpenStudy (anonymous):

Found value of y?

OpenStudy (usukidoll):

\[y=\frac{150}{13}\] 5y-6x=12 plug in y in this equation and solve for x

OpenStudy (usukidoll):

what are you talking about? I didn't post that.

OpenStudy (usukidoll):

\[5(\frac{150}{13})-6x=12\]

OpenStudy (anonymous):

@Kaelyn78 what you got for y? You break the interaction?? :(

OpenStudy (anonymous):

Okay, I gotta go now. If any doubt, Doll will tell you, or I will when I login next.. :) Solving these types of equations is really an easy task, just learn the concept, rest is a-jiffy work..

OpenStudy (anonymous):

sorry fell asleep

OpenStudy (anonymous):

I'm going to bed, thanks everyone, goodnight